A voltaic cell consists of a Pb/Pb2+ half-cell and a Cu/Cu2+ half-cell at 25 oC. The initial concentrations of Pb2+ and Cu2+ are 0.0500 M and 1.50 M, respectively. What are the concentrations of Pb2+ and Cu2+ when the cell potential falls to 0.370 V?

Cu2+(ag) + Pb(s) → Cu(s) + Pb2+(ag)       

Hint: [Pb2+] + [Cu2+] = (1.5M + 0.05M) = 1.55 M total

Hint: the standard potential of Pb2+ + 2e-   → Pb(s) is -0.130 and the standard potential of Cu2+ + 2e-  → Cu(s) is +0.340.

The equation is:

E= Eo- (RT/nF) (LnQ)
Answer #1

22 - Cell Reaction - cü ing) cute + Pbce) → (46) + P3+(ay) Escue = E catode - Eanode Ecattede = Einth/2y = 0.34V Eanode a EpsE= 0.37V 2- .. CV - CPT E - E 8.314 X218 2896480 Truth 0.01283 in - 50+] 70.05-x). 0.47 0.370 = In (1156+23 - 0.37V 0:47V -0.Final concertsation of U Crutz) = 0.05 cm = 0.08.-0.04g3 = 0.0007 Cpute] =oliste = 105 + 0.0493 = 1.5493